OpenStudy (anonymous):

how do you convert from kg/h^2 to g/s^2, and if you can explain the process please

OpenStudy (jamesj):

Well 1 kg = 1000 g , so 1 kg/h^2 = 1000 g/h^2 On the other hand 1 hr = 3600 seconds, so 1 1/h = 1/3600 1/s and thus 1 1/h^2 = 1/3600^2 1/s^2 Hence 1 kg/h^2 = 1000 g/h^2 = ...

OpenStudy (espex):

You can also use differential analysis with the basic units. So if you had 6kg/h^2 and you wanted to convert it to g/s^2 you would set it up like this: 1000g 1h^2 6kg ----- ------ ----- = 1kg 3600s^2 1h^2 Then cancel the units, multiply across the top and bottom, then divide the numbers and keep the remaining units.

OpenStudy (anonymous):

kk thanks you, how would i set up the analysis for g/cm to kg/m

OpenStudy (agent0smith):

1000g in 1kg 100cm in 1m So multiply the g/cm by 100 to convert to metres, then divide by 1000 to convert to kg. Or: \[\frac{ g }{ cm } \times \frac{ 100cm }{ 1m } \times \frac{ 1 kg }{1000g }\]And check that all the correct units cancel off, leaving you with only kg/m